# Deep SPI: Safe Policy Improvement via World Models

### v1

@@ 4 Your friendly neighborhood policy
Motivated by the intuition that constraining policy updates can mitigate OOT and confounding policy issues, we consider measuring the update as the importance ratio (IR) of the policies. This measure provides guarantees for constraining policy and representation updates, and with an appropriate optimisation scheme, ensures both policy improvement and convergence. In Section 5 we will further show that properly constraining the IR allows for safe policy improvements in world models while providing representation guarantees.
@@ 5 With great world models comes great representation
$L_{{R}}^{\mathcal{B}}\coloneqq\mathop{\mathbb{E}}_{s,a\sim\mathcal{B}}\left|{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\bar{s},a}}\right)\right|,$ $L_{P}^{\mathcal{B}}\coloneqq\mathop{\mathbb{E}}_{s,a\sim\mathcal{B}}\mathcal{W}\left(\phi_{\sharp}P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right),\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)\right)$
where $\phi_{\sharp}P$ is the pushforward measure of $P$ by $\phi$, and $\mathcal{W}$ the Wasserstein distance (Vaserstein, 1969). $\mathcal{W}$ between $\mu,\nu\in\Delta\mathopen{}\mathclose{{\left(\,\overline{\!{\mathcal{\mathcal{S}}}}}}\right)$ is defined as $\mathcal{W}\left(\mu,\nu\right)=\inf_{\lambda\in\Lambda\mathopen{}\mathclose{{\left(\mu,\nu}}\right)}\mathop{\mathbb{E}}_{\mathopen{}\mathclose{{\left(\bar{s},\bar{s}^{\prime}}}\right)\sim\lambda}\;\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\bar{s},\bar{s}^{\prime}}}\right)$, where $\Lambda(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$. While the Wasserstein operator may seem scary at first glance, it generalizes over transition losses that can be found in the literature. In particular, when the latent space is discrete, this distance boils down to the total variation distance. Another notable case is when the transition dynamics are deterministic, in which case the transition loss reduces to $L_{P}^{\mathcal{B}}=\mathop{\mathbb{E}}_{s,a,s^{\prime}\sim\mathcal{B}}\,\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}}\right)$. Finally, in general, a tractable upper bound can be obtained as $L^{\mathcal{B}}_{P}\leq\mathop{\mathbb{E}}_{s,a,s^{\prime}\sim\mathcal{B}}\mathop{\mathbb{E}}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}{\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\bar{s}^{\prime}}}\right)}$ (proof in Appendix C).
@@
For the sake of presentation, we restrict our attention to the episodic RL setting; we consider the standard RL framework where the environment is almost surely always eventually reset. Our results extend to general settings where a stationary distribution is accessible (cf. Remark LABEL:rmk:episodic-assumption, Appendix E).
World model quality. Before introducing our safe policy improvement theorem, we first show that the local losses effectively measure the world model’s quality with respect to the original environment. Namely, their difference in return obtained under any latent policy in a well-defined neighborhood is bounded by the local losses derived from the base policy’s state-action distribution. This is formalized in the following theorem.
@@
where $\textsc{AEL}\mathopen{}\mathclose{{\left(\pi_{\text{b}}}}\right)$ denotes the average episode length when $\mathcal{M}$ runs under $\pi_{\text{b}}$ and $K_{V}=\nicefrac{{K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\,\overline{\!{{R}}}}}}{{\mathopen{}\mathclose{{\left(1-\gamma K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}}}\right)}}$.
In simpler terms, if the deviation (supremum IR, or SIR for short) between the base policy and any new policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ stays stricly lower than $\nicefrac{{1}}{{\gamma}}$, the gap in return between the environment and the world model for this new policy can be bounded using data collected via $\pi_{\text{b}}$. Minimizing local losses from $\pi_{\text{b}}$’s data ensures that refining the representation $\phi$ for $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ improves model quality: when these losses vanish, $\mathcal{M}$ and $\overline{\!{\mathcal{M}}}$ are almost surely equivalent under $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$. The bound depends on the Average Episode Length (AEL), but even a loose upper bound is sufficient to preserve guarantees. It is also strongly influenced by the discount factor $\gamma$, which defines an implicit horizon. Smaller values permit larger deviations from $\pi_{\text{b}}$ and relax the accuracy required of the world model.
Safe policy improvement. We consider the setting where the world model is used to improve the baseline policy $\pi_{\text{b}}=\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$, with $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$ and the representation $\phi$ is fixed during each update. Restricting updates to a well-defined neighborhood guarantees that $\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi,\mathcal{M}}}\right)-\rho\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mathcal{M}}}\right)\geq\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{\mathcal{M}}}}}\right)-\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b},\,\overline{\!{\mathcal{M}}}}}\right)-\zeta$, where $\zeta$ is defined as the cumulative modeling error from the local losses.
@@
(Deep, Safe Policy Improvement) Under the same preamble as in Thm. 2, assume that $\phi$ if fixed during the policy update and the baseline is a latent policy with $\pi_{\text{b}}\coloneqq\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$ and $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$. Then, the improvement of the return of $\mathcal{M}$ under $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ can be guaranteed on $\pi_{\text{b}}$ as
@@
Theorem 3 addresses the OOT issue (Section 3.1): if the SIR of the baseline remains strictly below $\nicefrac{{1}}{{\gamma}}$, then minimizing the local losses reduces the error $\zeta$, ensuring safe policy improvement when the world model is used to enhance the policy. While our focus is not on offline SPI, Appendix E (Thm. LABEL:thm:deep-spi-pac) additionally provides a PAC variant of the result, following the standard use of confidence bounds in the SPI literature.
Representation learning. Finally, we analyze how learning a world model using our loss functions as an auxiliary task facilitates the learning of a useful representation. A good representation should ensure that environment states that are close in the representation also have close values, directly supporting policy learning. Specifically, we seek “almost” Lipschitz continuity (Vanderbei, 1991) of the form $\exists K\colon\forall s_{1},s_{2}\in\mathcal{S},\left|V^{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(s_{1}}}\right)-V^{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(s_{2}}}\right)\right|\leq K\cdot\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi_{\text{old}}\mathopen{}\mathclose{{\left(s_{1}}}\right),\phi_{\text{old}}\mathopen{}\mathclose{{\left(s_{2}}}\right)}}\right)+\mathcal{L}_{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(\phi_{\text{old}}}}\right)$ where $\mathcal{L}_{\pi_{\text{b}}}$ is an auxiliary loss depending on the data collected by $\bm{\pi_{\text{b}}}$. Notably, a critical question is whether updating the policy and its representation, respectively to $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ and $\phi$, maintains Lipschitz continuity. Crucially, as the baseline $\pi_{\text{b}}$ is updated to $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi$ with respect to the experience collected under $\pi_{\text{b}}$, the bound must hold for $\mathcal{L}_{\pi_{\text{b}}}$. The following theorem is a probabilistic version of this statement, formalized as a concentration inequality:
@@ 6 Across the SPI-verse: PPO comes into play
where $\ell_{R}\mathopen{}\mathclose{{\left(s,a}}\right)\coloneqq\left|{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)\right|$, $\ell_{P}\mathopen{}\mathclose{{\left(s,a,s^{\prime}}}\right)\coloneqq\mathop{\mathbb{E}}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}d\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\bar{s}^{\prime}}}\right)$, $s^{\prime}\sim P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right)$, and $\alpha_{R},\alpha_{P}\in\mathopen{(}0,1\mathclose{]}$. Intuitively, $\ell_{R}$, $\ell_{P}$ are transition-wise auxiliary losses that allow retrieving $L_{{R}}^{\xi_{\pi_{n}}}$ and $L_{P}^{\xi_{\pi_{n}}}$ in expectation w.r.t. the current policy $\pi_{n}$. When optimized, since they are clipped in a PPO-fashion, $U^{\pi_{n}}$ allows restricting the policy updates to the neighborhood.
@@
From this loss, we propose DeepSPI, a principled algorithm leveraging the policy improvement and representation learning capabilities developed in our theory. As our losses rely on distributions defined over the current policy, we focus on the on-policy setting. While model-based approaches are not standard in this setting, we stress that highly parallelized collection of data (e.g., via vectorized environments) enable a wide coverage of the state space (cf. Mayor et al., 2025; Gallici et al., 2025), which is suitable to optimize the latent model. DeepSPI updates the world model, the encoder, and the policy simultaneously while guaranteeing the representation is suited to perform safe policy updates.

### v2

@@ 4 Your friendly neighborhood policy
Motivated by the intuition that constraining policy updates can mitigate OOT and confounding policy issues, we consider measuring the update as the importance ratio (IR) of the policies. This measure provides guarantees for constraining policy and representation updates, and with an appropriate optimisation scheme, ensures both policy improvement and convergence. In Section 5, we will further show that properly constraining the IR allows for safe policy improvements in world models while providing representation guarantees.
@@ 5 With great world models comes great representation
$L_{{R}}^{\mathcal{B}}\coloneqq\mathop{\mathbb{E}}_{s,a\sim\mathcal{B}}\left|{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\bar{s},a}}\right)\right|,$ $L_{P}^{\mathcal{B}}\coloneqq\mathop{\mathbb{E}}_{s,a\sim\mathcal{B}}\mathcal{W}\left(\phi_{\sharp}P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right),\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)\right)$ (4)
where $\phi_{\sharp}P$ is the pushforward measure of $P$ by $\phi$, and $\mathcal{W}$ the Wasserstein distance (Vaserstein, 1969). $\mathcal{W}$ between $\mu,\nu\in\Delta\mathopen{}\mathclose{{\left(\,\overline{\!{\mathcal{\mathcal{S}}}}}}\right)$ is defined as $\mathcal{W}\left(\mu,\nu\right)=\inf_{\lambda\in\Lambda\mathopen{}\mathclose{{\left(\mu,\nu}}\right)}\mathop{\mathbb{E}}_{\mathopen{}\mathclose{{\left(\bar{s},\bar{s}^{\prime}}}\right)\sim\lambda}\;\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\bar{s},\bar{s}^{\prime}}}\right)$, where $\Lambda(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$. While the Wasserstein operator may seem scary at first glance, it generalizes over transition losses that can be found in the literature (cf. Sect. 1.1). In particular, when the latent space is discrete, this distance boils down to the total variation distance. Another notable case is when the transition dynamics are deterministic, in which case the transition loss reduces to $L_{P}^{\mathcal{B}}=\mathop{\mathbb{E}}_{s,a,s^{\prime}\sim\mathcal{B}}\,\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}}\right)$. Finally, in general, a tractable upper bound can be obtained as $L^{\mathcal{B}}_{P}\leq\mathop{\mathbb{E}}_{s,a,s^{\prime}\sim\mathcal{B}}\mathop{\mathbb{E}}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}{\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\bar{s}^{\prime}}}\right)}$ (proof in Appendix C).
@@
For the sake of presentation, we restrict our attention to the following assumption for Thms. 2 and 3:
Assumption 1.
We assume that the agent operates in the episodic RL setting, i.e., we consider the standard RL framework where the environment is eventually reset with probability one.
Our results extend to general settings where a stationary distribution is accessible (c.f. Remark LABEL:rmk:episodic-assumption).
World model quality. Before introducing our safe policy improvement theorem, we first show that the local losses effectively measure the world model’s quality with respect to the original environment. Namely, their difference in return obtained under any latent policy in a well-defined neighborhood is bounded by the local losses derived from the reference, behavioral policy’s state-action distribution. This is formalized in the following theorem.
@@
where $\textsc{AEL}\mathopen{}\mathclose{{\left(\pi_{\text{b}}}}\right)$ denotes the average episode length when $\mathcal{M}$ runs under $\pi_{\text{b}}$, $K_{V}=\nicefrac{{K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\,\overline{\!{{R}}}}}}{{\mathopen{}\mathclose{{\left(1-\gamma K^{\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu}_{\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu}}}\right)}}$, and $L_{{R}}^{\xi_{\pi_{\text{b}}}},L_{P}^{\xi_{\pi_{\text{b}}}}$ are the local losses of Eq. 4 over the stationary distribution $\xi_{\pi_{\text{b}}}$ induced by $\pi_{\text{b}}$.
In simpler terms, if the deviation (supremum IR, or SIR for short) between the behavioral policy and any new policy $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ stays strictly lower than $\nicefrac{{1}}{{\gamma}}$, the gap in return between the environment and the world model for this new policy can be bounded using data collected via $\pi_{\text{b}}$. Minimizing local losses from $\pi_{\text{b}}$’s data ensures that refining the representation $\phi$ for $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ improves model quality: when these losses vanish, $\mathcal{M}$ and $\overline{\!{\mathcal{M}}}$ are almost surely equivalent under $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$. The bound depends on the Average Episode Length (AEL), but even a loose upper bound is sufficient to preserve guarantees. It is also strongly influenced by the discount factor $\gamma$, which defines an implicit horizon. Smaller values permit larger deviations from $\pi_{\text{b}}$ and relax the accuracy required of the world model.
Safe policy improvement. We consider the setting where the world model is used to improve the behavioral policy $\pi_{\text{b}}=\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$, with $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$ and the representation $\phi$ is fixed during each update. Restricting updates to a well-defined neighborhood guarantees that $\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi,\mathcal{M}}}\right)-\rho\mathopen{}\mathclose{{\left(\pi_{\text{b}},\mathcal{M}}}\right)\geq\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu,\,\overline{\!{\mathcal{M}}}}}\right)-\rho\mathopen{}\mathclose{{\left(\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b},\,\overline{\!{\mathcal{M}}}}}\right)-\zeta$, where $\zeta$ is defined as the cumulative modeling error from the local losses.
@@
(Deep, Safe Policy Improvement) Under the same preamble as in Thm. 2, assume that $\phi$ if fixed during the policy update and the behavioral is a latent policy with $\pi_{\text{b}}\coloneqq\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\circ\phi$ and $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu_{b}\in\mkern 1.5mu\overline{\mkern-1.5mu\Pi\mkern-1.5mu}\mkern 1.5mu$. Then, the improvement of the return of $\mathcal{M}$ under $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ can be guaranteed on $\pi_{\text{b}}$ as
@@
Theorem 3 addresses the OOT issue (Section 3.1): if the SIR of the behavioral remains strictly below $\nicefrac{{1}}{{\gamma}}$, then minimizing the local losses reduces the error $\zeta$, ensuring safe policy improvement when the world model is used to enhance the policy. While our focus is not on offline SPI, Appendix E (Thm. LABEL:thm:deep-spi-pac) additionally provides a PAC variant of the result, following the standard use of confidence bounds in the SPI literature.
Representation learning. Finally, we analyze how learning a world model using our loss functions as an auxiliary task facilitates the learning of a useful representation. A good representation should ensure that environment states that are close in the representation also have close values, directly supporting policy learning. Specifically, we seek “almost” Lipschitz continuity (Vanderbei, 1991) of the form $\exists K\colon\forall s_{1},s_{2}\in\mathcal{S},\left|V^{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(s_{1}}}\right)-V^{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(s_{2}}}\right)\right|\leq K\cdot\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi_{\text{old}}\mathopen{}\mathclose{{\left(s_{1}}}\right),\phi_{\text{old}}\mathopen{}\mathclose{{\left(s_{2}}}\right)}}\right)+\mathcal{L}_{\pi_{\text{b}}}\mathopen{}\mathclose{{\left(\phi_{\text{old}}}}\right)$ where $\mathcal{L}_{\pi_{\text{b}}}$ is an auxiliary loss depending on the data collected by $\bm{\pi_{\text{b}}}$. Notably, a critical question is whether updating the policy and its representation, respectively to $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu$ and $\phi$, maintains Lipschitz continuity. Crucially, as the behavioral $\pi_{\text{b}}$ is updated to $\mkern 1.5mu\overline{\mkern-1.5mu\pi\mkern-1.5mu}\mkern 1.5mu\circ\phi$ with respect to the experience collected under $\pi_{\text{b}}$, the bound must hold for $\mathcal{L}_{\pi_{\text{b}}}$. The following theorem is a probabilistic version of this statement, formalized as a concentration inequality:
@@ 6 Across the SPI-verse: PPO comes into play
where $\ell_{R}\mathopen{}\mathclose{{\left(s,a}}\right)\coloneqq\left|{R}\mathopen{}\mathclose{{\left(s,a}}\right)-\,\overline{\!{{R}}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)\right|$, $\ell_{P}\mathopen{}\mathclose{{\left(s,a,s^{\prime}}}\right)\coloneqq\mathop{\mathbb{E}}_{\bar{s}^{\prime}\sim\mkern 1.5mu\overline{\mkern-1.5muP\mkern-1.5mu}\mkern 1.5mu\mathopen{}\mathclose{{\left({\cdotp}\mid\phi\mathopen{}\mathclose{{\left(s}}\right),a}}\right)}\,\overline{\!{d}}\mathopen{}\mathclose{{\left(\phi\mathopen{}\mathclose{{\left(s^{\prime}}}\right),\bar{s}^{\prime}}}\right)$, $s^{\prime}\sim P\mathopen{}\mathclose{{\left({\cdotp}\mid s,a}}\right)$, and $\alpha_{R},\alpha_{P}\in\mathopen{(}0,1\mathclose{]}$. Intuitively, $\ell_{R}$, $\ell_{P}$ are transition-wise auxiliary losses that allow retrieving $L_{{R}}^{\xi_{\pi_{n}}}$ and $L_{P}^{\xi_{\pi_{n}}}$ in expectation w.r.t. the current policy $\pi_{n}$. When optimized, since they are clipped in a PPO-fashion, $U^{\pi_{n}}$ allows restricting the policy updates to the neighborhood.
@@
From this loss, we propose DeepSPI, a principled algorithm leveraging the policy improvement and representation learning capabilities developed in our theory. As our losses rely on distributions defined over the current policy, we focus on the on-policy setting. While model-based approaches are not standard in this setting, we stress that highly parallelized collection of data (e.g., via vectorized environments) enables a wide coverage of the state space (cf. Mayor et al., 2025; Gallici et al., 2025), which is suitable to optimize the latent model. DeepSPI updates the world model, the encoder, and the policy simultaneously while guaranteeing the representation is suited to perform safe policy updates.
6.1 Illustrative Example
Figure 3: Toy maze environment illustrating the confounding policy update problem.
To illustrate the representation learning capabilities of DeepSPI, we consider the toy grid-world shown in Fig. 3. This environment mirrors the confounding policy update discussed in Sect. 3.2, instantiated earlier in Fig. 2.
The agent starts in the cell labeled I. Upon leaving the orange cell immediately to its right, it is sent to the top branch with probability $1-\epsilon$ and to the bottom branch with probability $\epsilon$. It must then traverse a corridor of $n$ blue cells (here nn=5). Moving one cell to the right yields a reward of +1, and the agent cannot move backwards.
At the final corridor cell, marked with a $\star$, moving right yields a reward of +1 regardless of whether the agent is in the top or bottom branch, and the episode terminates. The difference is when the agent moves up from the $\star$ cell: in the top branch, it receives a reward of $\nicefrac{{+n}}{{\gamma^{n}}}$, whereas in the bottom branch it receives $\nicefrac{{-(2-\epsilon)n}}{{\mathopen{}\mathclose{{\left(\epsilon\,\gamma^{n}}}\right)}}$ before termination. As in Sect. 3.2, this construction ensures that if both $\star$ states are merged in the latent space, choosing “right” remains acceptable (their values coincide), but choosing “up” produces a negative expected return from the initial state I (details in Appendix LABEL:appendix:toy-example). To improve upon the policy that always chooses “right,” the agent must learn to assign distinct representations to the two $\star$ cells, select “up” in the top one, and “right” in the bottom one.
We compare the behaviour of PPO and DeepSPI in this environment. Since our goal is to highlight the agent’s representation-learning capabilities, each observation is provided as raw pixels.
Figure 4: Value from cell I in the maze (left) and distance between the representation of the $\star$ cell from the top and bottom branches (right).
The agent must therefore learn both a policy and an encoder mapping pixels to a structured latent space. Further details on the environment and observation scheme are given in Appendix LABEL:appendix:toy-example. As shown in Fig. 4, the representation learned by PPO collapses the top and bottom $\star$ cells into a single latent state. With such a representation, the best policy PPO can learn is to always choose “right,” which leads to a return of $\sim\!4.8$. In contrast, DeepSPI benefits from the representation quality guarantees of Thm. 4, which ensure that states with different values remain separated in the latent space for all policies in a suitable neighborhood. This is exactly what we observe: the learned representation distinguishes the two $\star$ cells. As a result, the agent learns to choose “up” in the top $\star$ cell and “right” in the bottom one, achieving a return of $\sim\!8$.